# HydroGEM: A Self-Supervised Zero Shot Hybrid TCN-Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v2

@@ Abstract
Advances in sensor networks have enabled real-time, high-resolution stream discharge monitoring from in-situ gauges, yet persistent sensor malfunctions and data quality issues limit their utility. Manual quality control by expert hydrologists—while essential for maintaining data integrity—creates a bottleneck that delays scientific and operational use of these observations and cannot scale with networks generating millions of measurements annually. To address this gap, we introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a foundation model for continental-scale streamflow quality control designed to support human expertise rather than replace it. HydroGEM uses a two-stage training approach: self-supervised pretraining on 6.03 million clean sequences from 3,724 USGS stations learns general hydrological representations, followed by fine-tuning with synthetic anomalies for detection and reconstruction. A hybrid Temporal Convolutional Network–Transformer architecture (14.2M parameters) captures both local temporal patterns and long-range dependencies, while hierarchical normalization handles learning across six orders of magnitude in discharge. On a held-out test set of 799 stations with 18 expert-validated anomaly types, HydroGEM achieves F1 = 0.792 for detection and 68.7% reconstruction-error reduction, outperforming the strongest baseline by 36.3%. Zero-shot transfer to 100 Environment and Climate Change Canada stations—unseen during training—yields F1 = 0.586, demonstrating cross-national generalization. The model maintains consistent detection across correction magnitudes (1–100%) and aligns with operational seasonal patterns, with peak flagging rates during winter ice-affected periods matching hydrologists’ correction behavior. Architectural separation between simplified training anomalies and complex physical-space test anomalies confirms that performance reflects learned hydrometric principles rather than pattern memorization.
@@
Continental-scale foundation model trained on 3724 USGS sites with 6.03 million sequences
@@
Zero-shot transfer to Canadian ECCC stations shows cross-national generalization
@@ 1 Introduction
Advances in sensor technology, telemetry, and network infrastructure have dramatically expanded the spatial and temporal density of hydrological monitoring. Modern gauging stations transmit stage and discharge at sub-hourly intervals, enabling near-real-time flood warnings, reservoir operations, and ecological flow management. Yet despite these technological improvements, raw sensor data remain inherently imperfect. Even in well-maintained monitoring networks, physical sensors inevitably experience drift, fouling, ice effects, rating-curve shifts, clock errors, and transmission failures, producing intermittent gaps or anomalies in the data [51, 12]. These issues are an inherent part of real-world sensing, where the central challenge is ensuring fast, reliable quality control of the data they generate.
Traditional quality control in hydrological monitoring combines manual expert review with rule-based automated checks [4, 47, 24]. Commercial systems such as AQUARIUS implement configurable range checks, rate-of-change thresholds, and statistical comparison filters, providing standardized workflows and audit trails [4]. Open-source frameworks like SaQC advance reproducibility through sequential test configurations and traceable quality flags [47]. However, rule-based approaches face fundamental limitations: fixed thresholds require site-specific calibration and capture only obvious outliers. Context-dependent anomalies—such as ice effects that preserve plausible values but bias discharge [35], gradual sensor drift within historical ranges [53], or rating curve shifts after channel-altering floods [1]—often evade detection. Multivariate relationships between discharge and stage, critical for identifying coupled sensor failures or unit conversion errors, are difficult to encode in rule-based systems. As monitoring networks expand and temporal resolution increases, manual calibration becomes prohibitively expensive.
@@
Recent advances in machine learning offer a promising path toward more efficient and scalable hydrologic data quality control. Modern models can process continuous data streams, capture subtle temporal dependencies, and operate consistently across large station networks—capabilities well aligned with the needs of contemporary monitoring systems. In principle, such methods could provide rapid, reliable quality screening across thousands of sites.
@@
Deep learning approaches show stronger performance by capturing complex temporal dependencies. LSTM networks and Transformer attention mechanisms excel at modeling long-range patterns in time series [23, 21]. Temporal Convolutional Networks enable efficient parallel processing with exponentially large receptive fields [19]. Autoencoders—including variational and LSTM-augmented variants—detect anomalies via reconstruction errors [42, 31]. Generative Adversarial Networks improve robustness through learned representations of normal data distributions [8]. Synthetic anomaly generation has emerged as a key technique for training when domain-specific labeled data are unavailable [26].
@@
Foundation models—large networks pretrained on massive datasets with self-supervised objectives, then transferred to diverse downstream tasks—provide a path forward [10, 13]. In weather and climate science, ClimaX demonstrated that transformers pretrained on heterogeneous climate reanalysis data with masked token prediction effectively transfer to forecasting, downscaling, and multi-variable prediction across spatiotemporal scales [38]. Weather foundation models including FourCastNet [43], GraphCast [30], and Pangu-Weather [7] now achieve medium-range forecast skill competitive with operational numerical weather prediction at orders of magnitude lower computational cost. Industrial-scale efforts such as IBM/NASA’s Prithvi WxC (2.3B parameters) [48] and Oak Ridge’s ORBIT (>>100B parameters) [58] are extending this paradigm to multi-variable Earth system prediction.
@@
Yet foundation model principles have not been systematically applied to quality control for in situ sensor networks. This gap reflects unique challenges: extreme scale heterogeneity (six orders of magnitude in discharge within a single dataset, requiring specialized normalization); irregular multi-site time series (non-gridded data with site-specific rating curves, instrument types, and operational protocols); real-time deployment constraints (inference must operate without access to ground truth); deploy-safe requirements (the model must not corrupt valid data—a failure mode unacceptable in operational settings); and human oversight mandates (agencies require interpretable outputs, uncertainty quantification, and audit trails for regulatory compliance).
@@
HydroGEM uses a two-stage training approach. Stage 1 pretrains a hybrid TCN-Transformer backbone on 6.03 million clean sequences from 3,724 USGS sites using masked reconstruction. A hierarchical normalization scheme—combining log transforms, site-specific standardization, and explicit scale embeddings—allows learning across extreme magnitude ranges while preserving physically meaningful scale structure [17]. Stage 2 fine-tunes the pretrained backbone with a detection head using on-the-fly synthetic anomaly injection rather than curated labels [13, 26]. The training injector applies simplified corruptions (drift, spikes, flatlines, dropouts, clock shifts) in normalized space to encourage learning of fundamental hydrometric consistency principles rather than memorization of specific anomaly signatures. Clean data preservation exceeds 97%, ensuring the model does not corrupt valid observations.
Evaluation emphasizes generalization across three dimensions. We construct a synthetic test set of 799 USGS sites (completely unseen during training) with expert-validated anomalies that are deliberately more complex than training corruptions, creating a four-axis separation (geographic, mathematical, temporal, parameter) that prevents pattern memorization. On this held-out set, HydroGEM achieves 0.70 F1 for anomaly detection and reduces reconstruction error by 70% relative to injected corruptions—though suggested corrections require expert review [24] before operational use. For real-world validation, we evaluate zero-shot transfer to 100 Environment and Climate Change Canada (ECCC) sites using weak labels derived from operational corrections. Reconstructions show promising alignment with agency corrections in several cases, though direct comparison is limited by differences in correction methodology between USGS and ECCC workflows. This cross-national evaluation tests whether the learned representations transfer beyond the training data source [27, 11].
@@
    A continental-scale foundation model trained on 3,724 USGS sites with 6.03 million sequences—an order of magnitude larger than prior multi-site hydrological studies.
@@
    Expert-validated evaluation with 18 anomaly types and rigorous site-level partitioning ensuring zero-shot generalization assessment, achieving 0.70 F1 detection and 70% error reduction in suggested reconstructions.
@@
    Demonstrated cross-national transfer from USGS (USA) to ECCC (Canada) data, validating learned representations beyond training distribution.
@@
Table LABEL:tab:terminology summarizes key technical terms used in this paper.*** Ijaz, can we refer to the terminology table as a LaTeX table here?
Terminology Clean data: USGS approved records with quality code ‘A’ (verified and approved by USGS hydrologists as accurate representations of observed conditions; used for Stage 1 training) Masked data: Clean data with random timesteps hidden for self-supervised reconstruction (Stage 1) Corrupted data: Clean data with synthetic anomalies injected (Stage 2 fine-tuning) Anomalous data: Real operational data requiring correction (deployment target) Suggested reconstruction: Model-proposed correction requiring hydrologist approval Stage/Gage height: Water surface elevation measured at the gauging station, typically in feet Discharge: Volume of water flowing past a point per unit time, typically in cubic feet per second (ft3/s) Rating curve: Empirical relationship between stage and discharge at a specific site
@@ 2 Data and Evaluation Framework
Our evaluation strategy combines three data sources (Figure 1): (1) a continental-scale USGS corpus for foundation model pretraining, (2) synthetic test sets with expert-validated anomalies for controlled generalization assessment, and (3) real-world Canadian stations for zero-shot cross-national transfer validation. This multi-tiered design allows rigorous evaluation under both controlled conditions with known ground truth and operational settings with genuine quality control challenges.
@@ 2.1.1 USGS Streamflow Archive
We assembled a continental-scale dataset from 5,321 USGS streamgaging stations via the National Water Information System (NWIS) [55], covering January 2000 through December 2024. For each station, we selected the best 10-year continuous period with high-quality paired discharge ($Q$) and stage ($H$) observations. All records carry USGS qualification code ‘A’ (approved), distinguishing them from provisional data subject to revision. Instantaneous readings (typically 15-minute intervals) were aggregated to hourly resolution via arithmetic means for consistent temporal scale.
@@ 2.1.2 Spatial and Hydrologic Coverage
The network spans the conterminous United States, Alaska, and Hawaii (162.69°W to 67.32°W, 19.71°N to 61.71°N), providing true continental-scale geographic diversity. Drainage areas range from 0.03 to 932,800 km2 (median 196 km2), consistent with scale-invariant river network organization. Gauge elevations span 69-69 m to 2,747 m (mean 467 m), capturing coastal, lowland, montane, and alpine monitoring sites.
@@ 2.1.3 Quality Control and Preprocessing
Gap filling uses hierarchical strategies: linear interpolation for gaps $\leq 6$ hours, exponential recession models [57, 52] for 6-24 hour gaps, and exclusion for gaps >24>24 hours. This protocol retains 94.7% of potential sequences, with 89.3% containing no interpolation. Windows intersecting flagged spans are dropped, eliminating information leakage.
@@ 2.1.4 Dataset Partitioning
To assess generalization, we assign entire sites to a single partition, avoiding leakage of site-specific signatures [27]. The split allocates 70% training (3724 sites), 15% validation (798 sites), and 15% test (799 sites), preserving geographic coverage and proportional hydrologic region representation.
Continuous hourly series are segmented into 576-hour windows (24 days), capturing storm hydrographs, weekly patterns, and diurnal fluctuations while remaining tractable for transformers [56]. Asymmetric stride—48 hours for training (dense overlap maximizing data utilization), 192 hours for validation/test (reducing autocorrelation)—produces 6.03M training sequences, 0.32M validation, and 0.33M test sequences (Table 2).
@@ 2.2.1 Feature Engineering
At each timestep $t$ for site $s$, the model observes $\mathbf{x}_{t}\in\mathbb{R}^{12}$ combining static descriptors, dynamic hydrology, and derived temporal context (Table 3). This compact design balances information richness, computational tractability, and operational constraints—specifically, we exclude meteorological forcings (precipitation, temperature) to maximize applicability to gauge-only deployment scenarios.
@@
Scale embeddings $\sigma_{\ln Q}$, $\sigma_{\ln H}$ (training std. dev. of log-transformed series) 2
@@
Static descriptors $\{\phi,\lambda,A_{d},z_{g}\}$ provide coarse hydrograph controls through latitude (e.g., climatic gradients), longitude (e.g., continental position), drainage area (e.g., flashiness), and elevation (e.g., temperature lapse rates). Dynamic variables $\{Q,H\}$ constitute the core physical state, with stage encoding local geometry and backwater effects [39] invisible in discharge alone. Derived features include scale embeddings returning absolute variability information lost during standardization, cross-site ranks stabilizing learning across magnitude orders, and monthly anomalies expressing seasonal departures.

### v3

@@ Abstract
Advances in sensor networks have enabled real-time, high-resolution stream discharge monitoring from in-situ gauges, yet persistent sensor malfunctions and data quality issues limit their utility. Manual quality control by expert hydrologists creates a bottleneck that delays scientific and operational use of these observations and cannot scale with networks generating millions of measurements annually. To address this gap, we introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a foundation model for continental-scale streamflow quality control designed to support human expertise rather than replace it. HydroGEM uses a two-stage training approach: self-supervised pretraining on 6.03 million clean sequences from 3,724 USGS stations learns general hydrological representations, followed by fine-tuning with synthetic anomalies for detection and reconstruction. A hybrid Temporal Convolutional Network–Transformer architecture (14.2M parameters) captures both local temporal patterns and long-range dependencies, while hierarchical normalization handles learning across six orders of magnitude in discharge. On held-out real observations from 799 stations with synthetic anomalies spanning 18 types grounded in USGS operational standards, HydroGEM achieves F1 = 0.792 for detection and 68.7% reconstruction error reduction, outperforming the strongest baseline by 36.3%. For cross-national validation on 100 Environment and Climate Change Canada stations, we adopt tolerant evaluation with a $\pm$24 hour buffer to accommodate weak labels derived from operational corrections recorded with daily granularity. HydroGEM achieves Tolerant F1 = 0.70, with stable precision across all buffer sizes and 90.1% segment-level detection of anomaly events, demonstrating cross-national generalization. The model maintains consistent detection across correction magnitudes (1–100%) and aligns with operational seasonal patterns, with peak flagging rates during winter ice-affected periods matching hydrologists’ correction behavior. Architectural separation between simplified training anomalies and complex physical-space test anomalies confirms that performance reflects learned hydrometric principles rather than pattern memorization.
@@
Continental-scale foundation model trained on 3,724 USGS sites with 6.03 million sequences
@@
Zero-shot transfer to 100 Canadian ECCC stations achieves Tolerant F1 = 0.70
@@ 1 Introduction
Advances in sensor technology, telemetry, and network infrastructure have dramatically expanded the spatial and temporal density of hydrological monitoring. Modern gauging stations transmit stage and discharge at sub-hourly intervals, enabling real-time flood warnings, reservoir operations, and ecological flow management. Yet despite these technological improvements, raw sensor data remain inherently imperfect. Even in well-maintained monitoring networks, physical sensors inevitably experience drift, fouling, ice effects, rating-curve shifts, clock errors, and transmission failures, producing intermittent gaps or anomalies in the data [72, 14]. These issues are an inherent part of sensing in operational environments, where the central challenge is ensuring fast, reliable quality control of the data they generate.
Traditional quality control in hydrological monitoring combines manual expert review with rule-based automated checks [6, 65, 29]. Commercial systems such as AQUARIUS implement configurable range checks, rate-of-change thresholds, and statistical comparison filters, providing standardized workflows and audit trails [6]. Open-source frameworks like SaQC advance reproducibility through sequential test configurations and traceable quality flags [65]. However, rule-based approaches face fundamental limitations: fixed thresholds require site-specific calibration and capture only obvious outliers. Context-dependent anomalies often evade detection, including ice effects that preserve plausible values but bias discharge [47], gradual sensor drift within historical ranges [75], and rating-curve shifts after channel-altering floods [2]. Multivariate relationships between discharge and stage, critical for identifying coupled sensor failures or unit conversion errors, are difficult to encode in rule-based systems. As monitoring networks expand and temporal resolution increases, manual calibration becomes prohibitively expensive.
@@
Recent advances in machine learning offer a promising path toward more efficient and scalable hydrologic data quality control. Modern models can process continuous data streams, capture subtle temporal dependencies, and operate consistently across large station networks. These capabilities align well with the needs of contemporary monitoring systems. In principle, such methods could provide rapid, reliable quality screening across thousands of sites.
@@
Deep learning approaches show stronger performance by capturing complex temporal dependencies. LSTM networks and Transformer attention mechanisms excel at modeling long-range patterns in time series [25, 23]. Temporal Convolutional Networks enable efficient parallel processing with exponentially large receptive fields [21]. Hybrid designs that combine convolutional temporal feature extraction with attention based global context have also been explored for time series forecasting and anomaly detection [84, 1]. Autoencoders, including variational and LSTM-augmented variants, detect anomalies via reconstruction errors [55, 42]. Generative Adversarial Networks improve robustness through learned representations of normal data distributions [11]. Synthetic anomaly generation has emerged as a key technique for training when domain-specific labeled data are unavailable [32, 69].
@@
Foundation models provide a path forward [13, 15]. These large networks are pretrained on massive datasets with self-supervised objectives, then transferred to diverse downstream tasks. In weather and climate science, ClimaX demonstrated that transformers pretrained on heterogeneous climate reanalysis data with masked token prediction effectively transfer to forecasting, downscaling, and multi-variable prediction across spatiotemporal scales [51]. Weather foundation models including FourCastNet [37], GraphCast [39], and Pangu-Weather [10] now achieve medium-range forecast skill competitive with operational numerical weather prediction at orders of magnitude lower computational cost. Industrial-scale efforts such as IBM/NASA’s Prithvi WxC (2.3B parameters) [66] and Oak Ridge’s ORBIT (>>100B parameters) [83] are extending this paradigm to multi-variable Earth system prediction.
@@
Yet foundation model principles have not been systematically applied to quality control for in situ sensor networks. This gap reflects unique challenges: extreme scale heterogeneity (six orders of magnitude in discharge within a single dataset, requiring specialized normalization); irregular multi-site time series (non-gridded data with site-specific rating curves, instrument types, and operational protocols); real-time deployment constraints (inference must operate without access to ground truth); deploy-safe requirements (the model must not corrupt valid data, a failure mode unacceptable in operational settings); and human oversight mandates (agencies require interpretable outputs, uncertainty quantification, and audit trails for regulatory compliance). While hybrid convolution attention backbones are effective in general time series settings [84, 1], they typically do not address deploy-safe quality control requirements across thousands of heterogeneous gauging sites, including extreme magnitude variation, label scarcity, and cross-agency transfer constraints.
@@
HydroGEM uses a two-stage training approach. Stage 1 pretrains a hybrid TCN-Transformer backbone on 6.03 million clean sequences from 3,724 USGS sites using masked reconstruction. Unlike many prior hybrid architectures that use conventional softmax attention [84, 1], our global module adopts cosine attention with a retention style temporal decay prior to improve stability on long sequences [48, 73]. A hierarchical normalization scheme combines log transforms, site-specific standardization, and explicit scale embeddings to allow learning across extreme magnitude ranges while preserving physically meaningful scale structure [19]. Stage 2 fine-tunes the pretrained backbone with a detection head using on-the-fly synthetic anomaly injection rather than curated labels [15, 32]. The training injector applies simplified corruptions (drift, spikes, flatlines, dropouts, clock shifts) in normalized space to encourage learning of fundamental hydrometric consistency principles rather than memorization of specific anomaly signatures. Clean data preservation exceeds 97%, ensuring the model does not corrupt valid observations.
Evaluation emphasizes generalization across three dimensions. On held-out real observations from 799 USGS sites with synthetic anomalies spanning 18 types grounded in USGS operational standards, we create a four-axis separation (geographic, mathematical, temporal, parameter) that prevents pattern memorization. HydroGEM achieves F1 = 0.792 for anomaly detection and reduces reconstruction error by 68.7% relative to injected corruptions, outperforming the strongest baseline by 36.3%. Suggested corrections require expert review [29] before operational use. For cross-national validation, we evaluate zero-shot transfer to 100 Environment and Climate Change Canada (ECCC) sites using weak labels derived from operational corrections. Because correction records are often applied with daily granularity rather than precise anomaly boundaries, and correction methodologies vary across agencies [29], we adopt tolerant evaluation with a $\pm$24 hour buffer as the primary metric for this dataset. HydroGEM achieves Tolerant F1 = 0.70, with stable precision across all buffer sizes from $\pm$1 to $\pm$24 hours. At the segment level, the model detects 90.1% of anomaly events, demonstrating effective cross-national generalization despite differences in instrumentation and operational protocols.
@@
Minimal inference materials, runnable notebooks, and supporting documentation for the USGS synthetic benchmark and ECCC site set are available at https://huggingface.co/Ejokhan/HydroGEM.
@@
    A continental-scale foundation model trained on 3,724 USGS sites with 6.03 million sequences, an order of magnitude larger than prior multi-site hydrological studies.
@@
    Rigorous evaluation on held-out real observations from 799 sites with synthetic anomalies spanning 18 types grounded in USGS operational standards, achieving F1 = 0.792 for detection and 68.7% reconstruction error reduction.
@@
    Demonstrated cross-national transfer from USGS (USA) to ECCC (Canada) data with Tolerant F1 = 0.70, validating learned representations beyond the training distribution.
@@
Box 1 summarizes key technical terms used throughout this paper.
Box 1: Terminology
  • Clean data: USGS approved records with quality code ‘A’ (verified and approved by USGS hydrologists as accurate representations of observed conditions; used for Stage 1 training)
  • Masked data: Clean data with random timesteps hidden for self-supervised reconstruction (Stage 1)
  • Corrupted data: Clean data with synthetic anomalies injected (Stage 2 fine-tuning)
  • Anomalous data: Real operational data requiring correction (deployment target)
  • Suggested reconstruction: Model-proposed correction requiring hydrologist approval
  • Stage/Gage height: Water surface elevation measured at the gauging station, typically in feet
  • Discharge: Volume of water flowing past a point per unit time, typically in cubic feet per second (ft3/s)
  • Rating curve: Empirical relationship between stage and discharge at a specific site
@@ 2 Data and Evaluation Framework
Our evaluation strategy combines 3 data sources (Figure 1): (1) a continental-scale USGS corpus for foundation model pretraining, (2) synthetic test sets with anomalies grounded in USGS operational standards for controlled generalization assessment, and (3) real-world Canadian stations for zero-shot cross-national transfer validation. This design enables evaluation both under controlled conditions with known ground truth and under operational conditions that reflect agency quality control workflows.
@@ 2.1.1 USGS Streamflow Archive
We initially retrieved candidate records from 6,000 USGS streamgaging stations via the National Water Information System (NWIS) [79], covering January 2000 through December 2024. After applying the quality control and completeness filters described in Section 2.1.3, 679 stations were excluded, yielding 5,321 retained sites for analysis and partitioning. For each station, we selected the best 10-year continuous period with high-quality paired discharge ($Q$) and stage ($H$) observations. Best was defined as the continuous 10-year span maximizing paired $Q$ and $H$ completeness after quality control while minimizing gap filling. All retained records carry USGS qualification code ‘A’ (approved), distinguishing them from provisional data subject to revision. Instantaneous readings (typically 15-minute intervals) were aggregated to hourly resolution via arithmetic means for consistent temporal scale.
@@ 2.1.2 Spatial and Hydrologic Coverage
The network spans the conterminous United States, Alaska, and Hawaii, providing true continental-scale geographic diversity (approximately 163W to 67W and 20N to 62N). Drainage areas range from 0.03 to 932,800 km2 (median 196 km2). Gauge elevations span 69-69 m to 2,747 m (mean 467 m), capturing coastal, lowland, montane, and alpine monitoring sites.
@@ 2.1.3 Quality Control and Preprocessing
Gap filling uses hierarchical strategies: linear interpolation for gaps $\leq 6$ hours, exponential recession models [81, 74] for 6-to-24-hour gaps, and exclusion for gaps >24>24 hours. This protocol retains 94.7% of potential sequences, with 89.3% containing no interpolation. Windows intersecting flagged spans are dropped, eliminating information leakage.
@@ 2.1.4 Dataset Partitioning
To assess generalization, we assign entire sites to a single partition, avoiding leakage of site-specific signatures [34]. The split allocates 70% training (n = 3,724), 15% validation (n = 798), and 15% test (n = 799), preserving geographic coverage and proportional hydrologic region representation.
Continuous hourly series are segmented into 576-hour windows (24 days), capturing storm hydrographs, weekly patterns, and diurnal fluctuations while remaining tractable for transformers [80]. Asymmetric stride (48 hours for training to maximize data utilization through dense overlap, 192 hours for validation and test to reduce autocorrelation) produces 6.03M training sequences, 0.32M validation, and 0.33M test sequences (Table 2).
@@
The held-out test partition provides unseen sites for controlled anomaly benchmarking by injecting anomalies into these sites using the protocol described in Section 2.3.1.
@@ 2.2.1 Feature Engineering
At each timestep $t$ for site $s$, the model observes $\mathbf{x}_{t}\in\mathbb{R}^{12}$ combining static descriptors, dynamic hydrology, and derived temporal context (Table 3). This compact design balances information richness, computational tractability, and operational constraints. We intentionally exclude meteorological forcings (precipitation, temperature) to maximize applicability to gauge-only deployment scenarios.
@@
Scale embeddings $\sigma_{\ln Q}$, $\sigma_{\ln H}$ (training standard deviation of log-transformed series) 2
@@
Static descriptors $\{\phi,\lambda,A_{d},z_{g}\}$ provide coarse hydrograph controls through latitude (eg climatic gradients), longitude (eg continental position), drainage area (eg flashiness), and elevation (eg temperature lapse rates). Dynamic variables $\{Q,H\}$ constitute the core physical state, with stage encoding local geometry and backwater effects [52] not visible in discharge alone. Derived features include scale embeddings returning absolute variability information lost during standardization, cross-site ranks stabilizing learning across magnitude orders, and monthly anomalies expressing seasonal departures.